A simple pendulum makes 107 complete oscillations in 2.60 min at a location where g = 9.80 m/s2.

(a) Find the period of the pendulum.
s  

(b) Find the length of the pendulum.
m
Answer #1

Solution::

The frequency of the pendulum is f = 107sc/min = 2.60/s

The free fall acceleration g = 9.8 m/s^2

a)
Period of the pendulum is

T = 1/f = 1/2.60= 0.38 s

b)
The period os a simple pendulum is given by  

T= 2(pi) sqrt(L/g)

0.38= 2*3.14 *sqrt(L/9.8)

L = 0.036 m